### v1

Abstract
A Metropolis Monte Carlo algorithm is given for the case of a complex phase space weight, which applies generally in quantum statistical mechanics. Computer simulations using Lennard-Jones 4He near the $\lambda$-transition, including an expansion to third order of the Wigner-Kirkwood commutation function, give a saturation liquid density in agreement with measured values.
I Introduction
The computer simulation of quantum condensed matter is a challenging problem. A review of methods then extant has been given by Ceperley (1995). His own quantum Monte Carlo algorithm that implements the Feynman path integral over imaginary time (equivalently, temperature slices) is the most impressive. Ceperley (1995) illustrates his algorithm with extensive results for saturated liquid 4He in the vicinity of the $\lambda$-transition.
One significant development since 1995 is the exact formulation of quantum statistical mechanics in classical phase space (Attard 2016b, 2018b, 2021). This has been used for Monte Carlo simulations of the noble gases (Attard 2017), with more recent focus on liquid helium-4 in the vicinity of the $\lambda$-transition (Attard 2025a, 2025b, 2025d). Arguably, Ceperley’s (1995) Feynman path integral algorithm does a better job at handling the non-commutativity of the position and momentum operators (also known as the Wigner-Kirkwood commutation function, also known as the Heisenberg uncertainty relation), whereas Attard’s classical phase space algorithm is better at handling wave function symmetrization.
The computational burdens of the two approaches are very different. Ceperley (1995) simulated 64 4He atoms using a 1990’s supercomputer. Attard (2025a, 2025b, 2025d) simulated 5,000 4He atoms using a naughty personal computer. The primitive approximation for the action on the Feynman path requires about 1,000 temperature slices (Ceperley 1995), the cost of each one of which is comparable to a single classical simulation. Ceperley (1995) gives more sophisticated approximations for the action that require far fewer slices, ${\cal O}(20)$, but the additional computational cost of each slice is unclear.
The purpose of the present paper is to improve the classical phase space algorithm with respect to the Wigner-Kirkwood commutation function, whilst preserving its advantages for the treatment of wave function symmetrization, and without increasing the computational burden. A general Metropolis Monte Carlo algorithm is given for complex phase space weights. This is implemented for Lennard-Jones 4He on the saturation curve. A fourth order temperature expansion of the Wigner-Kirkwood commutation function is given, and numerical results are obtained up to the third order. Methods for including the symmetrization function are given, but these are not implemented in the present paper.
II Formalism and Algorithm
II.1 Phase Space Weight
For a subsystem of $N$ identical bosons in a volume $V$, a point in classical phase space is ${\bf\Gamma}=\{{\bf q},{\bf p}\}$, and its conjugate with reversed momenta is ${\bf\Gamma}^{\dagger}=\{{\bf q},-{\bf p}\}$. Here the position configuration is ${\bf q}=\{{\bf q}_{1},{\bf q}_{2},\ldots,{\bf q}_{N}\}$ and the momentum configuration is ${\bf p}=\{{\bf p}_{1},{\bf p}_{2},\ldots,{\bf p}_{N}\}$, where the position of boson $j$ is ${\bf q}_{j}=\{{q}_{jx},{q}_{jy},{q}_{jz}\}$, and its momentum is ${\bf p}_{j}=\{{p}_{jx},{p}_{jy},{p}_{jz}\}$. Here we take the momentum to belong to the continuum; see Attard (2025a, 2025b) for the treatment of quantized momentum.
For a canonical equilibrium system of temperature $T$ and volume $V$, the phase space probability density is (Attard 2018b, 2021)
$\wp({\bf\Gamma})=\frac{e^{-\beta{\cal H}({\bf\Gamma})}e^{W({\bf\Gamma})}\eta({\bf\Gamma})}{N!h^{3N}Z},$ (2.1)
where $\beta=1/k_{\rm B}T$, $k_{\rm B}$ being Boltzmann’s constant, $h$ is Planck’s constant, and the partition function, $Z(N,V,T)$, which normalizes the probability, gives the total entropy. Also the classical Hamiltonian is ${\cal H}({\bf\Gamma})={\cal K}({\bf p})+U({\bf q})$, where ${\cal K}({\bf p})=p^{2}/2m=\sum_{j=1}^{N}p_{j}^{2}/2m$ is the kinetic energy, $m$ being the mass of a boson. Below we shall take the potential energy to consist of central pair potentials $U({\bf q})=\sum_{j<k}^{N}u(q_{jk})$.
II.1.1 Commutation Function
The main focus of this work is on the Wigner-Kirkwood commutation function (Attard 2018b, 2021, Kirkwood 1933, Wigner 1932), which is defined by
$e^{-\beta{\cal H}({\bf\Gamma})}e^{W({\bf\Gamma})}=e^{{\bf p}\cdot{\bf q}/{\rm i}\hbar}e^{-\beta\hat{\cal H}({\bf q})}e^{-{\bf p}\cdot{\bf q}/{\rm i}\hbar}.$ (2.2)
This corresponds to $\omega_{p}=e^{W_{p}}$ (Attard 2018b Eq. (2.6)) (present notation). For the quantum weight one can use either $\omega_{p}\eta_{q}$ or else $\omega_{q}\eta_{p}$, with $\omega_{q}^{*}=\omega_{p}$ and $\eta_{q}^{*}=\eta_{p}$.
On the right hand side of the definition of the commutation function the Hamiltonian operator appears, $\hat{\cal H}({\bf q})=\hat{\cal K}({\bf q})+U({\bf q})$. The Fourier factors are unnormalized, unsymmetrized momentum eigenfunctions. The commutator $\hat{\cal K},U]\neq 0$ makes $W\neq 0$, which reflects the Heisenberg uncertainty relation.
The Wigner-Kirkwood commutation function is complex with the property that $W({\bf\Gamma}^{\dagger})=W({\bf\Gamma})^{*}$. Separating the real and imaginary parts, $W=W_{\rm r}+{\rm i}W_{\rm i}$, we have
$e^{W({\bf\Gamma})}=e^{W_{\rm r}({\bf\Gamma})}[\cos W_{\rm i}({\bf\Gamma})+{\rm i}\sin W_{\rm i}({\bf\Gamma})].$ (2.3)
The imaginary part of this is odd in momentum and averages to zero because in an equilibrium system forward and backward momentum must be equally likely, $\wp({\bf\Gamma})=\wp({\bf\Gamma}^{\dagger})$. Thus we only require the real part to obtain averages of real functions. For the moment we neglect the fact that the symmetrization function $\eta$ is complex. The commutation function becomes non-zero at higher temperatures than the symmetrization function, and so we can deduce intrinsic properties of the former in the absence of the latter.
The cosine term rapidly fluctuates about zero, particularly since $W$ is an extensive thermodynamic variable. Such rapid oscillations cancel each other, and the corresponding region of phase space has zero weight. Therefore, we impose the restriction
$|W_{\rm i}({\bf\Gamma})|<\frac{\pi}{2}.$ (2.4)
The justification for this is that at high temperatures the system is classical and $W({\bf\Gamma})=0$. Presumably, as the temperature is lowered the system never crosses the boundary where $\cos W_{\rm i}({\bf\Gamma})$ changes sign. As will shortly be shown, the Wigner-Kirkwood commutation function depends upon the gradients of the pair potential. Since these become large in the repulsive core region of the pair potential, this restriction prevents particles approaching each other too closely. (Some gradients appear as the scalar product with the momentum, and so this also tends to lower the magnitude of the latter.) As the temperature is lowered, higher order gradients come into play, and these are large at greater separations than lower order gradients. This restriction is a manifestation of the Heisenberg uncertainty relation: it keeps particles further apart than classical considerations alone imply. This condition is perhaps best judged by the results that follow.
A standard approach to the simulation of classical equilibrium systems is Metropolis Monte Carlo (Allen and Tildesley 1987). In the most general form this says that a trial move to a new configuration should be accepted if $\wp^{\rm new}/\wp^{\rm old}\geq r$, where $r$ is a random number uniformly distributed on $[0,1]$. In the present case this says accept the trial configuration if
$e^{-\beta\Delta{\cal H}({\bf\Gamma})}e^{\Delta W_{\rm r}({\bf\Gamma})}\frac{\cos W_{\rm i}({\bf\Gamma}^{\rm new})}{\cos W_{\rm i}({\bf\Gamma}^{\rm old})}\geq r.$ (2.5)
This is the major result of this paper. Note that both position and momentum are changed in a trial move, which is generally done one particle at a time. In addition, the trial configuration is rejected if the restriction (2.4) is violated. Hence if we start with $\wp>0$, then this means that the probability always remains positive.
[…]
II.1.2 Symmetrization Function
In the computational results for 4He presented below the symmetrization function is neglected. This is appropriate on the high temperature side of the $\lambda$-transition. Our main aim is to delineate the rôle of the commutation function, and we shall take up the combined effects of the two in future work. For completeness, here we show one way in which this can be done.
The symmetrization function is the ratio of unpermuted to permuted momentum eigenfunctions summed over all permutations (Attard 2018b, 2021),
$\eta({\bf\Gamma})=\sum_{\hat{\rm P}}e^{-[{\bf p}-\hat{\rm P}{\bf p}]\cdot{\bf q}/{\rm i}\hbar}.$ (2.6)
This corresponds to $\eta_{q}$ (Attard 2018b Eq. (2.4)). That part of the grand potential due to symmetrization is given by the average of this,
$e^{-\beta\Omega_{\rm sym}}$ $=$ $\langle\eta({\bf\Gamma})\rangle_{W}$ (2.7)
$=$ $e^{\langle\stackrel{{\scriptstyle\circ}}{{\eta}}({\bf\Gamma})\rangle_{W}}.$
In the second equality, which is believed to be exact in the thermodynamic limit (Attard 2018b § III B 1), the sum over single permutation loops is $\stackrel{{\scriptstyle\circ}}{{\eta}}\!\!({\bf\Gamma})=\sum_{l=2}^{\infty}\eta^{(l)}({\bf\Gamma})$, where the $l$-loop symmetrization function is
$\eta^{(l)}({\bf\Gamma})=\sum_{j_{1},\ldots,j_{l}}^{N}\!\!\!{}^{\prime}\;\prod_{k=1}^{l}e^{-{\bf p}_{j_{k}}\cdot{\bf q}_{j_{k},j_{k+1}}/{\rm i}\hbar},\quad j_{l+1}\equiv j_{1}.$ (2.8)
The sum is over the unique directed cyclic permutations of all subsets of $l$-bosons.
The $l$-loop symmetrization function is complex, $\eta^{(l)}({\bf\Gamma})=\eta^{(l)}_{\rm r}({\bf\Gamma})+{\rm i}\eta^{(l)}_{\rm i}({\bf\Gamma})$. When averaged it gives the $l$-loop grand potential. Again because the imaginary part is odd in momentum, only the real part of the integrand survives, $\mbox{Re}\{e^{{\rm i}W_{\rm i}(\Gamma)}\eta^{(l)}(\Gamma)\}=\cos\big(W_{\rm i}(\Gamma)\big)\eta^{(l)}_{\rm r}({\bf\Gamma})-\sin\big(W_{\rm i}(\Gamma)\big)\eta^{(l)}_{\rm i}({\bf\Gamma})$. (This is why it is important to combine $W_{p}$ with $\eta_{q}$, or else $W_{q}$ with $\eta_{p}$.) Hence the $l$-loop grand potential is (Attard 2018b, 2021)
$-\beta\Omega_{W}^{(l)}$ $=$ $\langle\eta^{(l)}(\Gamma)\rangle_{W}$
$=$ $\left\langle\big[\eta^{(l)}_{\rm r}({\bf\Gamma})-\tan\!\big(W_{\rm i}(\Gamma)\big)\eta^{(l)}_{\rm i}({\bf\Gamma})\big]\right\rangle_{W}$
$=$ $\frac{\int{\rm d}{\bf\Gamma}\,e^{-\beta{\cal H}}e^{W_{\rm r}}\big[\eta^{(l)}_{\rm r}\cos W_{\rm i}-\eta^{(l)}_{\rm i}\sin W_{\rm i}\big]}{\int{\rm d}{\bf\Gamma}\,e^{-\beta{\cal H}}e^{W_{\rm r}}\cos W_{\rm i}}.$
(The monomer grand potential is just the logarithm of the partition function in the absence of the symmetrization function.) Several tricks have been found to facilitate the computation of the $l$-loop symmetrization function (Attard 2021 §5.4.2). These will undoubtedly prove useful in the present more complex case of numerical momentum quadrature. On the high temperature side of the $\lambda$-transition, we expect that only position loops with consecutive particles separated by less than about the thermal wavelength will contribute.
II.2 Fluctuation Expansion for the Commutation Function
Kirkwood (1933) took the inverse temperature derivative of the defining equation for the commutation function, Eq. (2.2), to obtain an infinite series in powers of $\beta$, explicitly giving the first non-zero term, which is quadratic in $\beta$. Higher order coefficients using this method have been obtained (Attard 2021 §§8.2–8.4).
A slightly different approach based on a fluctuation expansion has $n$th order coefficient (Attard 2021 §8.5)
$\Delta_{\cal H}^{(n)}({\bf q},{\bf p})\equiv\frac{\langle{\bf q}|\big[\hat{\cal H}-{\cal H}({\bf q},{\bf p})\big]^{n}|{\bf p}\rangle}{\langle{\bf q}|{\bf p}\rangle}.$ (2.10)
One has $\Delta^{(0)}_{\cal H}({\bf q},{\bf p})=1$ and $\Delta^{(1)}_{\cal H}({\bf q},{\bf p})=0$. These give a series for the exponent of the form
$W(\Gamma)=\frac{\beta^{2}}{2!}\tilde{\Delta}_{\cal H}^{(2)}(\Gamma)-\frac{\beta^{3}}{3!}\tilde{\Delta}_{\cal H}^{(3)}(\Gamma)+\frac{\beta^{4}}{4!}\tilde{\Delta}_{\cal H}^{(4)}(\Gamma)$-$\ldots$ (2.11)
The second order fluctuation, $\tilde{\Delta}^{(2)}_{\cal H}=\Delta^{(2)}_{\cal H}$, is
$\Delta^{(2)}_{\cal H}({\bf q},{\bf p})$ $=$ $\frac{-\hbar^{2}}{2m}\nabla^{2}U({\bf q})-\frac{\mathrm{i}\hbar}{m}{\bf p}\cdot\nabla U({\bf q}).$ (2.12)
The third order fluctuation, $\tilde{\Delta}^{(3)}_{\cal H}=\Delta^{(3)}_{\cal H}$, is
$\Delta^{(3)}_{\cal H}({\bf q},{\bf p})$ $=$ $\frac{-\hbar^{2}}{m}\nabla U\!\cdot\!\nabla U+\frac{\hbar^{4}}{4m^{2}}\nabla^{2}\nabla^{2}U$
$\mbox{ }+\frac{\mathrm{i}\hbar^{3}}{m^{2}}{\bf p}\!\cdot\!\nabla\nabla^{2}U-\frac{\hbar^{2}}{m^{2}}{\bf p}{\bf p}:\nabla\nabla U.$
And the fourth order fluctuation is
$\tilde{\Delta}^{(4)}_{\cal H}({\bf q},{\bf p})$ $=$ $\frac{5\hbar^{4}}{2m^{2}}\nabla U\cdot\nabla\nabla^{2}U+\frac{5\mathrm{i}\hbar^{3}}{m^{2}}{\bf p}\nabla U:\nabla\nabla U$ (2.14)
$\mbox{ }+\frac{\hbar^{4}}{m^{2}}\nabla\nabla U\!:\!\nabla\nabla U-\frac{\hbar^{6}}{8m^{3}}\nabla^{2}\nabla^{2}\nabla^{2}U$
$\mbox{ }-\frac{3\mathrm{i}\hbar^{5}}{4m^{3}}{\bf p}\!\cdot\!\nabla\nabla^{2}\nabla^{2}U+\frac{3\hbar^{4}}{2m^{3}}{\bf p}{\bf p}:\nabla\nabla\nabla^{2}U$
$\mbox{ }+\frac{\mathrm{i}\hbar^{3}}{m^{3}}{\bf p}{\bf p}{\bf p}\vdots\nabla\nabla\nabla U.$
Expressions for the gradients of a central pair potential have been catalogued (Attard 2021 §§9.5.2 and 9.5.3).
The order at which the expansion is terminated is denoted $n_{W}^{\rm max}$. The value $n_{W}^{\rm max}=0$ corresponds to a classical simulation.
II.3 Temperature Derivatives
The inverse temperature derivative of the logarithm of the partition function, without $\eta$, is
$\tilde{E} \displaystyle\equiv$ $\frac{-\partial\ln Z}{\partial\beta}$ (2.15)
$=$ $\frac{-1}{2Z}\int{\rm d}{\bf\Gamma}\,\left\{e^{-\beta{\cal H}}e^{W_{\rm r}}e^{{\rm i}W_{\rm i}}\left[\dot{W}_{\rm r}-{\cal H}+{\rm i}\dot{W}_{\rm i}\right]\right$.
 $\left.\mbox{ }+e^{-\beta{\cal H}}e^{W_{\rm r}}e^{-{\rm i}W_{\rm i}}\left[\dot{W}_{\rm r}-{\cal H}-{\rm i}\dot{W}_{\rm i}\right]\right\}$
$=$ $\frac{-1}{Z}\int{\rm d}{\bf\Gamma}\,e^{-\beta{\cal H}}e^{W_{\rm r}}\left\{[\dot{W}_{\rm r}-{\cal H}]\cos W_{\rm i}-\dot{W}_{\rm i}\sin W_{\rm i}\right\}$
$=$ $\left\langle{\cal H}({\bf\Gamma})-\dot{W}_{\rm r}({\bf\Gamma})+\dot{W}_{\rm i}({\bf\Gamma})\tan W_{\rm i}({\bf\Gamma})\right\rangle_{W}.$
The over-dot signifies the inverse temperature derivative, $\partial/\partial\beta$. The second derivative is
$\frac{\partial\tilde{E}}{\partial\beta}$ $=$ $\frac{(\partial Z/\partial\beta)^{2}}{Z^{2}}-\frac{1}{2Z}\int{\rm d}{\bf\Gamma}\,e^{-\beta{\cal H}({\bf\Gamma})}e^{W_{\rm r}({\bf\Gamma})}e^{{\rm i}W_{\rm i}({\bf\Gamma})}$ (2.16)
$\mbox{ }\times\left\{[-{\cal H}+\dot{W}_{\rm r}+{\rm i}\dot{W}_{\rm i}]^{2}+[\ddot{W}_{\rm r}+{\rm i}\ddot{W}_{\rm i}]\right\}+{\rm cc}$
$=$ $\left\langle{\cal H}-\dot{W}_{\rm r}+\dot{W}_{\rm i}\tan W_{\rm i}\right\rangle^{2}_{W}$
$\mbox{ }-\left\langle{\cal H}^{2}+\dot{W}_{\rm r}^{2}-\dot{W}_{\rm i}^{2}-2{\cal H}\dot{W}_{\rm r}+\ddot{W}_{\rm r}\right$.
 $\left.\mbox{ }-[-2{\cal H}\dot{W}_{\rm i}+2\dot{W}_{\rm r}\dot{W}_{\rm i}+\ddot{W}_{\rm i}]\tan W_{\rm i}\right\rangle_{W}.$
Note that the heat capacity is $C_{V}=\partial\tilde{E}/\partial T=-k_{\mathrm{B}}\beta^{2}\partial\tilde{E}/\partial\beta$.
III Computational Results
III.1 Model and Simulation Details
The Lennard-Jones pair potential was used, $u(r)=4\varepsilon[(\sigma/r)^{12}-(\sigma/r)^{6}]$, with helium parameters, $\varepsilon_{\rm He}=10.22k_{\rm B}$ J and $\sigma_{\rm He}=0.2556$ nm (van Sciver 2012). The potential was cut-off at $R_{\rm cut}=3.5\sigma$. Quantum Monte Carlo simulations were performed in classical phase space with either 1,000 or 5,000 atoms. A low total density, $\rho\sigma^{3}\approx$ 0.02–0.2, allowed a droplet of density $\rho\sigma^{3}\approx$ 0.3–0.9 to condense in the center of the system surrounded by its vapor. Periodic boundary conditions were applied. A small cell spatially based neighbor table was used. The position and momentum steps for a trial configuration were chosen to give acceptance rates of 30–60%.
As the temperature is decreased, the higher order gradients of the potential in the expansion for the commutation function come into play. These become very large at smaller separations. It was found that occasionally pairs of particles became trapped in the core region, presumably due to numerical overflow. The problem more commonly arose when changing density or temperature, and using the previous equilibrated configuration to start. The effect could be seen in the form of noise in the otherwise zero core region of the radial distribution function; this had amplitude ${\cal O}(10^{-3})$, compared to its peak height of ${\cal O}(10^{0})$. Although this has negligible effect on the statistics, there is no need to calculate zero explicitly. Therefore a hard core was implemented, with value $q_{\rm min}=$ 1.1–1.4, depending on the temperature. Any trial move that bought an atom closer to a neighbor than this was rejected. It was checked at the end of each simulation that the radial distribution went smoothly to zero and that the first peak was well-separated from the imposed hard-core.
For the fourth order case, $n_{W}^{\rm max}=4$, the commutation function became very large, ${\cal O}(10^{4})$, when any two particles came close together, $q_{jk}\lesssim 1.6\sigma$. This was due to the terms with a product of the gradients of the pair potential. Unfortunately the change in the exponent of the probability ratio was similarly large and led to numerical overflow. The problem could probably have been overcome by using a larger minimum separation $q_{\rm min}$, or a smaller position step to obtain the trial position configuration. However the former was deemed uncomfortably large, and the latter would have meant an unacceptably lengthy simulation. In so far as the fourth order term is a fluctuation of a fluctuation, (Attard 2021 Eq. (8.98)), it and higher order terms may well be negligible. No results for $n_{W}^{\rm max}=4$ are reported below.
III.2 Results
Table 1: Simulation results for the density $\rho\sigma^{3}$ and kinetic energy $\beta{\cal K}/N$ of liquid 4He on the saturation curve using the Wigner-Kirkwood commutation function with $n_{W}^{\rm max}=$0, 2, or 3 ($N=1,000$).
$k_{\mathrm{B}}T/\varepsilon$ $\Lambda/\sigma$ 0a 2 3 $\rho_{\rm meas}^{\rm sat}\sigma^{3}$ $\beta{\cal K}/N$b
0.8 1.1940 0.8023 0.30(1) 0.1912(6) - 1.1878(5)
0.7 1.2764 0.8470 0.398(2) 0.2027(6) - 1.1505(6)
0.6 1.3787 0.8872 0.446(2) 0.229(1) - 1.0930(4)
0.5 1.5103 0.9331 0.483(4) 0.2743(5) 0.255c 0.975(1)
0.45 1.5920 0.96 - 0.3236(5) 0.289 0.728(1)
0.45d - - - 0.3216(3) - 0.705(1)
0.4 1.6886 0.98 - 0.34(2) 0.317 0.642(5)
0.35 1.8051 - - 0.32(1) 0.339 0.580(2)
0.3 1.9498 - - 0.32(7) 0.355 0.501(1)
0.25 2.1359 - - 0.3324(1) 0.364 0.454(3)
[…]
Simulation results are shown in Table 1. For the classical simulation, $n_{W}^{\rm max}=0$, the saturated liquid density is about three times the measured value. Including the commutation function at second order reduces the density by about a factor of two, and including it to third order reduces it to about the measured value.
The kinetic energy per particle is $3k_{\rm B}T/2$ in the classical case and also for $n_{W}^{\rm max}=2$. In the third order case the kinetic energy is reduced from this value, increasingly so as the temperature is decreased.
A number of simulations were carried out with $N=5,000$, and the results of one such is shown in the table. No size dependence was observed. Similarly, simulations were repeated with different over-all densities without observing any systematic effects.
The measured critical temperature of 4He in Lennard-Jones units is $k_{\mathrm{B}}T^{\rm crit}_{\rm meas}/\varepsilon=0.48$. Table 1 gives results for the liquid saturation density above this temperature. In these cases a liquid droplet in a gas phase could be discriminated in the system. Evidently the critical temperature for the Lennard-Jones model is higher than the measured critical temperature, and it depends on the order of the commutation function expansion.
At $T^{*}=0.5$ the effective energy is $\beta\tilde{E}/N=-10.15(1)$, and the heat capacity is $C_{V}/Nk_{\rm B}=39.6(19)$. At $T^{*}=0.45$ these are $\beta\tilde{E}/N=-19.49(2)$ and $C_{V}/Nk_{\rm B}=59.7(23)$.
Figure 1: Radial distribution function for Lennard-Jones 4He at $k_{\rm B}T/\varepsilon=0.5$ at the respective liquid saturation densities. The dotted curve is $n_{W}^{\rm max}=0$ (classical, $\rho\sigma^{3}=0.9331$), the dashed curve is $n_{W}^{\rm max}=2$ ($\rho\sigma^{3}=0.483(4)$, $q_{\rm min}=1.1\sigma$), and the dotted curve is $n_{W}^{\rm max}=3$ ($\rho\sigma^{3}=0.2790(2)$, $q_{\rm min}=1.3\sigma$).
Figure 1 shows the radial distribution function for different orders of the commutation function. The statistics were collected over the whole system, but they are dominated by the saturated liquid phase. It can be seen that in the second and third order cases, the peak of the distribution occurs at a significantly greater separation than the imposed hard-core diameter, and that the distribution goes smoothly to zero. The signature of using too large a value for the hard-core is that the distribution goes discontinuously to zero immediately from the peak.
Notice how the peak of the radial distribution function shifts to larger separations as the order of the expansion for the commutation function is increased. This is a manifestation of the Heisenberg uncertainty relation, which smears out or delocalises bound particles. This is the origin of the reduction in the saturated liquid density compared to the classical case.
Figure 2: Density profile along an axis through the system at $k_{\rm B}T/\varepsilon=0.5$ (dashed curve) and at $k_{\rm B}T/\varepsilon=0.45$ (solid curve) for $n_{W}^{\rm max}=3$. The statistical error (95% confidence) is about 4% for the former, and about 20% for the latter.
Figure 2 shows the density in cross-section of the system at two temperatures. The density was averaged in a cylinder of radius $\sigma$ about an axis. It can be seen that at $k_{\rm B}T/\varepsilon=0.5$ the droplet is liquid-like, and that at $k_{\rm B}T/\varepsilon=0.45$ it is solid-like. In the latter case there is still substantial diffusion of the atoms over the course of the simulation. This and the fact that the troughs don’t go to zero means that it is not a rigidly crystalline state. Of course, even a liquid displays density oscillations at the liquid-vapor interface. The reason for describing the state as solid-like is the qualitative change in density profiles over such a small temperature change. As the temperature is further reduced, the peaks become more pronounced, the troughs can go to zero, and the diffusion becomes more limited. The density quoted in Table 1 is the average over a sphere of radius $8\sigma$ about the origin.
Real 4He does not become solid at saturation pressure down to absolute zero, but it does solidify at elevated pressures. One might speculate that the present solid-like state results from a combination of two effects. The first is the approximate nature of the Lennard-Jones pair potential together with the termination of the temperature expansion for the commutation function at $n_{W}^{\rm max}=3$. The second is the Laplace pressure acting on the interior of the nano-droplet.
IV Conclusion
A benefit of the present third order approximation for the Wigner-Kirkwood commutation function is that allows the classical phase space simulation of Lennard-Jones 4He at about the measured saturation liquid density in a homogeneous system. In the absence of the commutation function such a system would cavitate into a liquid drop at the bare Lennard-Jones density, $\rho_{LJ}^{\rm sat}\sigma^{3}\approx 0.9$, and a vapor phase. Being able to use the measured saturation liquid density for a homogeneous simulation is a significant advance.

### v2

@@
 $\mbox{ }-\left\langle{\cal H}^{2}+\dot{W}_{\rm r}^{2}-\dot{W}_{\rm i}^{2}-2{\cal H}\dot{W}_{\rm r}+\ddot{W}_{\rm r}\right$.
- $\left.\mbox{ }-[-2{\cal H}\dot{W}_{\rm i}+2\dot{W}_{\rm r}\dot{W}_{\rm i}+\ddot{W}_{\rm i}]\tan W_{\rm i}\right\rangle_{W}.$
-Note that the heat capacity is $C_{V}=\partial\tilde{E}/\partial T=-k_{\mathrm{B}}\beta^{2}\partial\tilde{E}/\partial\beta$.
+ $\left.\mbox{ }-[-2{\cal H}\dot{W}_{\rm i}+2\dot{W}_{\rm r}\dot{W}_{\rm i}+\ddot{W}_{\rm i}]\tan W_{\rm i}\right$.
+ $\left.\mbox{ }-\dot{W}_{\rm i}^{2}\sec^{2}W_{\rm i}\right\rangle_{W}.$
+(The final term was missing in version 1.) Note that the heat capacity is $C_{V}=\partial\tilde{E}/\partial T=-k_{\mathrm{B}}\beta^{2}\partial\tilde{E}/\partial\beta$.
 III Computational Results
@@
 Evidently the critical temperature for the Lennard-Jones model is higher than the measured critical temperature, and it depends on the order of the commutation function expansion.
-At $T^{*}=0.5$ the effective energy is $\beta\tilde{E}/N=-10.15(1)$, and the heat capacity is $C_{V}/Nk_{\rm B}=39.6(19)$.
-At $T^{*}=0.45$ these are $\beta\tilde{E}/N=-19.49(2)$ and $C_{V}/Nk_{\rm B}=59.7(23)$.
+For $k_{\rm B}T/\varepsilon=0.50$ and $\rho_{\rm tot}\sigma^{3}=0.26$ (droplet, central density 0.290(9), edge oscillations), gives $\beta\overline{E}/N=-12.41(4)$ and $\beta^{2}\partial\overline{E}/N\partial\beta=34(2)$.
+For $k_{\rm B}T/\varepsilon=0.49$, $\beta\overline{E}/N=-13.41(4)$ and $\beta^{2}\partial\overline{E}/N\partial\beta=37(1)$.
+(These results have been recalculated from version 1.)
+The numerical difference is $\Delta E/Nk_{\rm B}\Delta T=37.06(3)$.
+That this agrees with the directly simulated heat capacity is a very sensitive test of the mathematical expressions for the energy and the heat capacity, and of their computer implementation.
+It also tests the Metropolis Monte Carlo algorithm and its implementation.
+The author can confirm that it takes many rounds of debugging to get quantitative agreement between the two.
+The energy and heat capacity are sensitive to the saturated droplet and vary with the total density.
+For example, at $k_{\rm B}T/\varepsilon=0.50$, an overall density $\rho_{\rm tot}\sigma^{3}=0.15$.
+forms a smooth droplet with central density 0.28 and gives $\beta\overline{E}/N=-10.37(1)$.
+An overall density $\rho_{\rm tot}\sigma^{3}=0.03$ forms a smooth droplet with central density 0.28 and gives $\beta\overline{E}/N=-8.6(1)$.
+In these the energy of all atoms in the system are included.
 Figure 1: Radial distribution function for Lennard-Jones 4He at $k_{\rm B}T/\varepsilon=0.5$ at the respective liquid saturation densities.
@@
 Real 4He does not become solid at saturation pressure down to absolute zero, but it does solidify at elevated pressures.
-One might speculate that the present solid-like state results from a combination of two effects.
+One might speculate that the present solid-like state results from a combination of three effects.
 The first is the approximate nature of the Lennard-Jones pair potential together with the termination of the temperature expansion for the commutation function at $n_{W}^{\rm max}=3$.
 The second is the Laplace pressure acting on the interior of the nano-droplet.
+The third is the periodic boundary conditions.
 IV Conclusion